### v1

# Multi-Agents are Social Groups: Investigating Social Influence of Multiple Agents in Human-Agent Interactions

Abstract.
Multi-agent systems - systems with multiple independent AI agents working together to achieve a common goal - are becoming increasingly prevalent in daily life. Drawing inspiration from the phenomenon of human group social influence, we investigate whether a group of AI agents can create social pressure on users to agree with them, potentially changing their stance on a topic. We conducted a study in which participants discussed social issues with either a single or multiple AI agents, and where the agents either agreed or disagreed with the user’s stance on the topic. We found that conversing with multiple agents (holding conversation content constant) increased the social pressure felt by participants, and caused a greater shift in opinion towards the agents’ stances on each topic. Our study shows the potential advantages of multi-agent systems over single-agent platforms in causing opinion change. We discuss design implications for possible multi-agent systems that promote social good, as well as the potential for malicious actors to use these systems to manipulate public opinion.
Social Influence, Multi-agent, LLM Agent, Persuasion
*Equal contribution
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 18–22, 2025; XXX, XXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Psychologyccs: Human-centered computing User studies
Figure 1. Overview of our study.
1. Introduction
AI agents are advanced systems capable of interacting with users, holding conversations, and expressing their own thoughts (Wang et al., 2024b). These agents leverage the conversational and reasoning capabilities of large language models (LLMs) like Gemini, GPT-4, and LLaMA, and have become commonplace in daily life (Reuters, 2024). Significant research has also explored using these agents across a broad range of scenarios including education (Lieb and Goel, 2024), search interfaces (Ma et al., 2024), and trading platforms (Yu et al., 2024). The prevalence and ubiquity of LLM-powered agents has given rise to the concept of multi-agent systems - systems with two or more independent agents working together to fulfill a common goal and improve task performance (Guo et al., 2024; Chen et al., 2023b; Clarke et al., 2022). Major AI companies such as Anthropic (Claude, 2024) and Google (Google, 2024) have recently introduced multi-agent frameworks, making it easier for developers to create such interfaces.
With both the rise of multi-agent systems and the increasing prevalence of virtual agents in our daily lives, an interesting question emerges: Can a group of AI agents collectively exert social influence on users? This arises from a well-documented observation in human-human interaction: individuals are more likely to be influenced by the opinions of larger groups of people, a phenomenon known as social influence. Various forms of social influence, including conformance to the majority (Cialdini and Goldstein, 2004), peer pressure (Shepherd et al., 2011; Hui and Buchegger, 2009), desire for social acceptance (Richters and Waters, 1991; Oostveen et al., 1996), and adherence to social norms (McDonald and Crandall, 2015; Goldstein and Cialdini, 2011), are pervasive in everyday life. It is well-known that human-computer interaction dynamics tend to mimic human-human dynamics (Nass et al., 1994), with humans perceiving and reacting to computer systems similar to how they would behave towards other humans. This leads to the possibility that a group of AI agents holding the same opinion may exert social influence on users, similar to how a group of humans would.
The importance of investigating this issue is two-fold. Firstly, multi-agent interactions have become increasingly common. Most notably, on social media platforms like Twitter, human-like social bots have been effectively used to disseminate information and increase user engagement (Wischnewski et al., 2024). Multi-agent systems have also been designed to support online discussions (Zhang et al., 2024), public information elicitation (Jiang et al., 2023), and health coaching (Beinema et al., 2021), exposing users to opinions from multiple agents in a range of settings. Secondly, the potential social influence of multi-agent systems on users raises potential concerns. Evidence from psychology studies suggests that interactions with groups sharing similar views can significantly impact user opinions (Myers and Lamm, 1976; Isenberg, 1986), potentially leading to opinion polarization and manipulation. As digital agents increasingly function as social actors, the influence they exert on users is a critical yet underexplored issue.
Beyond potential negative effects, the social influence of multi-agent systems can also be harnessed in a positive way. In the traditional human-human domain, majority opinion has long been used to encourage changes in attitudes and behaviors, such as in health interventions (Skalski and Tamborini, 2007; Zhang et al., 2015) and sustainability efforts (Athanasiadis and Mitkas, 2005; Vossen et al., 2009). In human-computer interaction, persuasive systems have been designed across numerous domains like health and wellness (Balloccu et al., 2021; Oyebode et al., 2021), education (Ahtinen and Kaipainen, 2020; Tian et al., 2021), and social welfare (Shi et al., 2020), aiming to improve user behavior or change their attitudes through persuasion and informational education. Yet, these implementations are limited to single-agent interactions. The potential of multi-agent systems that collectively effect changes in user behavior or opinion remains unexplored.
The potential of AI agents in enacting persuasive effects is further supported by research in the HCI domain, which has highlighted the significant social influence individual AI agents exert, shaping users’ attitudes (Jakesch et al., 2023) and behaviors (Shi et al., 2020; Zhang et al., 2020). However, such research has focused on the design characteristics of individual agents: for instance, on how the perceived identity of an agent as a chatbot or a human impacts outcomes (Shi et al., 2020). Little attention has been given to how the number of agents affects the effectiveness of these systems. Existing research on multi-agent system has largely revolved around enhancing task performance (Guo et al., 2024; Chen et al., 2023b), either by helping users complete tasks more effectively or improving the technical performance of the system itself (Du et al., 2023; Chen et al., 2023a). In contrast, the indirect effects of social influence caused by these multi-agent systems has been largely overlooked. The impact of social influence is less relevant in these systems, where the primary goal is often to improve the clarity of information conveyed to the user. In scenarios where the system needs to both convey information and persuade a user of a certain viewpoint, the effect of multiple vs. single agents is critically underexplored.
Therefore, given 1) the increasing prevalence of multi-agent systems and interactions with multiple social bots online, 2) the known social influence that AI agents can exert, and 3) the established impacts of group social influence among humans that may be mirrored in human-AI interactions, there is 4) significant potential for both positive and negative effects caused by the collective social influence created by a group of AI agents. In this study, we address this issue by asking: Can interacting with multiple virtual agents create social influence that shifts our opinions? Specifically, we explore whether interactions with multiple agents can induce greater levels of social influence and lead to stronger opinion changes, and identify demographics factors that increase susceptibility to AI influence, and aim to form guidelines for the effective and ethical implementation of multi-agent systems.
To investigate these questions, we conducted an experiment (n = 94) where participants discussed their stances on two social topics with one, three, or five agents. Each participant was randomly assigned to a condition, with discussion content kept consistent across groups, differing only in the number of agents presenting it. Agents presented arguments that aligned with the participant’s stance on one topic and opposed it on the other. Survey data and open-ended responses were collected before and after each conversation round.
Through quantitative and qualitative analysis, we found that the multi-agent setup significantly influenced opinion shifts. Participants showed stronger opinion changes toward the agents when they disagreed and greater polarization when agents agreed with them. However, the shift did not intensify as the number of agents increased from 3 to 5. Instead, the 5-agent group led to a sense of polarization, with more participants rejecting the agents’ arguments and moving further from the agents’ positions. Additionally, multi-agent groups heightened perceptions of social influence, particularly normative influence to align with the majority. Open-ended responses indicated that participants expressed a desire for affiliation within the group, which contributed to social influence and opinion shifts in the 3- and 5-agent conditions. Demographic analysis further revealed that younger participants were more susceptible to multi-agent influence.
Our study makes the following contributions:
  • We show that having multiple agents in a discussion is more likely to create social influence and drive opinion change. This addresses a gap in understanding how multi-agent systems influence people, revealing their potential to drive stronger shifts in attitudes compared to single-agent setups.
  • We provide a foundational understanding of the differences in social influence between single-agent and multi-agent systems. By observing participants’ feelings of social exclusion and their desire to engage, we identify an implicit sense of affiliation with a group of agents, expanding existing social influence theories to include human-agent interactions.
  • We provide design suggestions for using multi-agent setups to more effectively convey persuasive strategies. By leveraging the normative influence exerted by multiple agents, future systems could foster emotional engagement in participants, promoting more lasting and internalized changes in attitudes and behaviors.
2. Related Work
2.1. Multi-Agent Interactions
AI agents are increasingly integrated into our daily life. These advanced AI systems are capable of interacting with users, holding conversations, and expressing their own thoughts (Wang et al., 2024b). Recent research and industry applications have leveraged the conversational and reasoning capabilities of LLMs to create problem-solving tools (Wang et al., 2024a; Roy et al., 2024; Ning et al., 2024; Yu et al., 2024; Guan et al., 2024), educational assistants (Lieb and Goel, 2024), and novel search interfaces (Ma et al., 2024). In the industry, many companies deploy AI agents on social media to assist with branding and customer service (Xu et al., 2017; Hu et al., 2018; Jiang et al., 2022), while others embed agents in their applications for assisting writing (Dhillon et al., 2024), brainstorming (Rezwana and Maher, 2023), or gaming (Sidji et al., 2024).
Despite this, there are many tasks that a single LLM-based agent struggles to handle effectively. Individual LLMs can lack sufficient domain expertise to handle specialized tasks (Ge et al., 2023), and may have insufficiently robust reasoning capabilities. Inspired by existing psychological theories (Smoliar, 1991), studies have investigated techniques such as ”debating” (Du et al., 2023) or ”reconciling” (Chen et al., 2023a) between multiple agents, to enhance reasoning by leveraging diverse perspectives. This has driven growing interest in ”multi-agent collaboration,” where agents work together to improve task performance (Guo et al., 2024; Chen et al., 2023b). Fields such as natural language processing (Guo et al., 2024), software engineering (Cardoso and Ferrando, 2021), and robotics (Chen et al., 2024) have leveraged the collective intelligence of multiple agents to simulate a group dynamic and improve task performance (Qian et al., 2023; Hong et al., 2023; Xiong et al., 2023). Multi-agent setups have also been used to simulate different human roles, akin to multiple workers interacting (Park et al., 2022, 2023b) or collaborating on a task (Light et al., 2023; Aher et al., 2023).
In more user-centered contexts, there have been emerging efforts to enable interactions with multiple agents simultaneously. For example, Clarke et al. (Clarke et al., 2022) developed an interface to facilitate engagement with multiple conversational agents at once, combining the unique features of different smart assistants like Google Assistant, SoundHound, and Ford’s in-vehicle smart assistant Adasa (Clarke et al., 2022). Additionally, many major AI companies like Anthropic (Claude, 2024), Google (Google, 2024), and others (Coze, 2024) are introducing multi-agent frameworks that enable developers to easily deploy multi-agent interfaces. Some organizations also use a multi-bot setup on social media to interact with and engage users.
Given the potential prevalence of human and multi-agent collaboration in everyday life, the interaction between multiple agents and humans remains underexplored in the HCI domain. Previous work has a limited focus on multi-agent systems in HCI, mostly highlighting opportunities of multi-agent interfaces in improving user outcomes: either by reducing cognitive load in information gathering (Jiang et al., 2023) or to provide information from diverse sources to improve decision making (Tan and Liew, 2022; Park et al., 2023a; Beinema et al., 2021; Chaves and Gerosa, 2018). These studies also show the challenges that multi-agent interfaces create for users, such as causing users to feel overwhelmed or confused by multiple agents providing different inputs (Chaves and Gerosa, 2018; Tan and Liew, 2022; Park et al., 2023a). In contrast, no existing work to our knowledge has investigated if and how a group of agents affects users through social influence mechanisms. Such effects are distinct from the cognitive and decision making effects in previous studies, yet are potentially as important.
In this study, we investigate the scenario: what if multiple agents are perceived as social actors, and they all share the same opinion on a topic just like a group of people might? This situation can arise, for instance, when companies deploy multiple agents on social media with a unified stance. While the influence of a single opinionated bot is well-documented (Jakesch et al., 2023; Shi et al., 2020; Zhang et al., 2020), the impact of being surrounded by multiple like-minded bots is less understood. Considering the growing prevalence of multi-agent systems, it is crucial to investigate how they might shape people’s opinions. Drawing from existing theories on human behavior within social groups, we hypothesize several potential effects, which we discuss in the next section.
2.2. Social Influence Theory
When multiple people express the same opinion, others in the group often feel compelled to agree. This phenomenon is known as ”social influence” (Cialdini and Goldstein, 2004), which comprises the ways in which individuals adjust their behavior to meet the demands of a social environment. This theory is well-known and widely studied in social psychology, and has been applied across different contexts, such as marketing (Salganik et al., 2006; Zhu et al., 2012; Teo et al., 2019), health intervention (Skalski and Tamborini, 2007; Zhang et al., 2015), sustainability (Athanasiadis and Mitkas, 2005; Vossen et al., 2009) and political discussions (Price et al., 2006).
Social influence is a broad term that manifests in various forms. Two main types of social influence are normative influence, the desire to ”conform with the expectations of another”, and informational influence, the desire to ”accept information obtained from another as evidence about reality” (Deutsch and Gerard, 1955). Normative influence relates to many common types of social influence including conformity (Cialdini and Goldstein, 2004), peer pressure (Shepherd et al., 2011; Hui and Buchegger, 2009), socialization (Richters and Waters, 1991; Oostveen et al., 1996), and social norms (McDonald and Crandall, 2015; Goldstein and Cialdini, 2011). More recent literature has termed this mechanism affiliation (Cialdini and Goldstein, 2004), or a desire to fit in with the majority: assimilation into a social group is often desired, and agreeing with the majority opinion can expedite this process. Naturally, normative social influence is more common among individuals that form a collective social group (Deutsch and Gerard, 1955), which creates an environment that one might want to be a part of. In this study, we investigate whether a group of AI agents can create a similar sense of a social group, implicitly encouraging a user to join the group by way of agreeing with their opinions.
Informational influence, on the other hand, relates to the concept of accuracy (Cialdini and Goldstein, 2004), or the natural instinct to gravitate towards objectively correct opinions as expressed by a group. This is distinct from normative influence as individuals may ”accept an opponent’s beliefs as evidence about reality even though one has no [other] motivation to agree with him” (Deutsch and Gerard, 1955). Despite this, social influence can still persuade individuals to accept an opinion, as the perceived consensus (Cialdini and Goldstein, 2004) of a larger number of people expressing the same viewpoint may suggest that that viewpoint is accurate. One distinction between human-human and human-AI interactions is that in human groups, each person represents a distinct source of knowledge and possible truth; with AI agents, it is trivial to set up multiple agents based on the same source of data, meaning that the number of agents presenting an opinion has no relation to its objective accuracy. Nevertheless, the tendency of humans to perceive such agents as distinct social actors (Nass et al., 1994) raises the possibility of informational influence from a group of agents.
In HCI research, a handful of studies (e.g. (Adam et al., 2021)) have examined the social influence that individual chatbots can exert on humans. In this study, we aim to conduct, to our knowledge, the first investigation of the possibility of both normative and informational social influence caused by groups of AI agents.
2.3. Conversational Agents in Online Discussions
Conversational agents (or ”chatbots”) powered by LLMs such as GPT-4 are capable of holding human-like conversations, and are increasingly being used in group discussion settings. Chatbots have long been used in online discussion scenarios, such as to debate with participants and prompt critical thinking (Tanprasert et al., 2024) or facilitate group discussion by improving discussion efficiency and members’ participation (Hadfi et al., 2023; Kim et al., 2020, 2021). Building on these conversational abilities, chatbots have also been used for persuasive means, to encourage a user or a group of users to alter their attitudes or behaviors. Such systems have been used in health and wellness (Balloccu et al., 2021; Oyebode et al., 2021), education (Ahtinen and Kaipainen, 2020; Tian et al., 2021) and social welfare (Shi et al., 2020). While these systems have demonstrated successful persuasion of their users, researchers have also discussed how design characteristics of the chatbot - such as the perceived identity of the bot (i.e., chatbot or human) (Shi et al., 2020) - can affect its persuasiveness.
Given the widespread application of chatbots for engaging users in real-world scenarios and the growing research interest in this area, the impact of the number of chatbots on persuasive outcomes — specifically the use of multi-agents in persuasion — remains underexplored. Leveraging the potential social influence effects of multi-agent systems may offer a novel means for designers of persuasive systems to enhance their effectiveness: for instance, health and wellness systems (Balloccu et al., 2021; Oyebode et al., 2021) could be designed with multiple agents rather than a single one, to increase the persuasiveness of the intervention. With the increasing accessibility of multi-agent platforms (Section 2.1), designers across various fields are more likely to encounter and develop such systems, making the exploration of their resulting social effects a critical area of investigation.
Given this, we raise the question of whether chatbots and LLM agents are capable of exerting social influence in the same ways that human users do. In particular, we investigate whether the affiliation and accuracy mechanisms work with virtual agents. HCI research regarding the Computers as Social Actors (CASA) phenomenon (Nass et al., 1994) has shown various ways in which humans perceive and act towards machines as though they were human. We extend this line of work by asking: do users feel inclined to agree with a group of virtual agents due to a desire to fit in with them?
Specifically, we propose the following research questions:
RQ1: Do interactions with multi-agent systems lead to stronger opinion changes?
This question aims to capture the direct outcomes of human interaction with multiple agents. Specifically, we hypothesize that when agents agree with users, users’ opinions will become more polarized in multi-agent conditions. Conversely, when agents disagree with users, their opinions will shift closer to the agents’ stance. We address this research question using both quantitative and qualitative methods.
RQ2: Do interactions with multi-agent systems lead to stronger social influence from agents?
To answer this question, we quantitatively test the hypothesis that as the number of agents increases, users will experience stronger social influence. We also use qualitative analysis of open-ended responses to explore the underlying reasons for this social influence. Here, we aim to identify potential mechanisms that drive perceived social influence in interactions with multiple agents and to draw comparisons with mechanisms found in human social group interactions.
RQ3: Which user demographics are more likely to be influenced by multi-agent systems?
Lastly, we explore whether participants with different demographic backgrounds are more easily influenced by multi-agent systems, by analyzing the relationship between demographic factors and the outcomes in RQ1 and RQ2 using regression analysis.
3. Methods
To understand how single- and multi-agent systems create social influence on humans, we conducted a mixed-methods study comprising an experiment and multiple survey segments, utilizing both quantitative measures and qualitative open-ended questions. This study received approval from our university’s ethics review committee prior to commencement.
Figure 2. Overview of the experiment. Part (a) illustrates the three experimental conditions—1-agent, 3-agent, and 5-agent. Part (b) displays example screenshots of each condition during the self-introduction phase. Part (c) outlines the study procedure, starting with a pre-survey, followed by two rounds of discussions (each followed by a mid-survey), and concluding with a post-survey.
3.1. Experiment Setup
To investigate how the number of agents influences human opinions, we randomly assigned participants to interact with one, three, or five agents. We chose these numbers based on prior research on persuasion and multi-agent interface designs: on one hand, these numbers represent different group sizes in human communication (Cacioppo and Petty, 1979; Trost et al., 1992); on the other, they reflect a practical range of agent numbers for real-world multi-agent applications (Beinema et al., 2021; Jiang et al., 2023; Park et al., 2023a).
To enhance the generalizability of our findings, we adopted two measures: (1) We chose two social topics for discussion instead of one. Specifically, drawing from recent HCI research on social discussions, we chose the topics “Should self-driving cars be allowed on roads?” (Govers et al., 2024) and “Do violent video games promote violence in youths?” (Yeo et al., 2024). We selected these two topics because they are highly relevant to the everyday lives of our participants, making it easier for them to have opinions and thoughts on them. (2) We designed our agent(s) to have two different stances — agents agreeing with participants on one topic and disagreeing on another — to examine whether alignment or disagreement with participants’ opinions affects the degree of social influence and opinion shift.
The study procedure is shown in Figure 2. Participants first completed a pre-survey that assessed their initial opinions on the social topics and their prior experience related to these topics. They were then randomly assigned to one of three conditions: 1-agent, 3-agent, or 5-agent. In each condition, a host agent appeared to guide the participants through the study. The host agent first introduced the task, then asked all agents and the participant to give a brief self-introduction. Then, for each topic, participants engaged in two rounds of conversation with their assigned agent(s). In each conversation round, the host agent first introduced the topic, then asked the agent(s) or participants to share their opinions on it. After the participants shared their viewpoints, the agent(s) responded to the participants’ statements of the topics and express their own thoughts. After each round of conversations, participants completed a mid-experiment survey to rate their perceived social influence of the agent(s) and their perceptions of the interactions. Once all conversations were finished, participants completed a post-survey to capture their final opinions on the two topics.
Figure 3. Example dialog showing the same set of arguments presented by different numbers of agents.
3.2. Agent Setup
The agents’ conversations were implemented through a combination of rule-based scripts and the GPT-4 API111gpt-4-1106-preview; https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4. For rule-based scripts, we designed a series of arguments either supporting or opposing the stance for each topic, as detailed in Appendix A.3. These arguments were then crafted into agent dialogue, such as ”I would definitely support [self-driving cars] because I think self-driving cars are great! You can sit back, relax, and let the car do the work for you” or ”I couldn’t agree with [allowing self-driving cars] because I think there are still a lot of issues with self-driving cars that need to be addressed. For example, technical developments are not yet perfect.” Across the three conditions, the same set of arguments was presented: with three and five agents, the different agents took turns to present different arguments; with one agent, the same agent presented all the arguments in the same order (see Figure 3 for an example of the dialogue). This was to ensure that any difference in opinion shifts across the three conditions was not due to differences in content quality or quantity.
We also integrated GPT-4 to enhance agent conversations in two ways: (1) parsing user input, such as extracting the user’s name from their greeting, and (2) generating interactive responses during discussions. When the host agent prompted users to share their opinions on the topics, the agent(s) provided brief feedback based on the user’s inputs, such as by giving a summary of the user’s thoughts or asking follow-up questions. For example, if a user said, ”I agree that self-driving cars should be allowed on the roads”, the next agent might have responded with, ”That’s great to hear! What aspects of self-driving cars do you find most appealing or beneficial?” These responses were tightly controlled through pre-set prompts (Appendix A.4), ensuring that the agent(s) only gave responses relevant to the discussion.
The agents’ avatars and rhetorical styles were designed to be human-like to enhance user acceptance (Sheehan et al., 2020). To avoid the uncanny valley effect (Song and Shin, 2024), we used cartoon-style avatars instead of realistic photos. Each avatar was assigned unique colors to help participants distinguish between the agents. We also ensured gender balance within the 3-agent and 5-agent conditions to minimize the potential effect of agent gender on participants’ opinion changes (Tanprasert et al., 2024).
Figure 4. Screenshot of user interface. The conversation page features an instruction at the top, guiding the user on how to begin the study. Users can enter text in the input field located at the bottom. Once the conversation begins, it is directed by a host agent who provides ongoing guidance throughout.
3.3. Participants
Participants were recruited via CloudResearch [^2] https://www.cloudresearch.com/. The selection criteria required them to be English speakers and over 18 years old. A total of 104 participants were initially recruited, with two not completing the survey and eight failing the attention check. Ultimately, 94 participants were included in the analysis: 31 participants (F: 17, M: 14) in the 1-agent group, 33 participants (F: 17, M: 15) in the 3-agent group, and 30 participants (F: 15, M: 15) in the 5-agent group. The average ages for each group were as follows: 1-agent group = 38.48 (SD = 12.41), 3-agent group = 38.75 (SD = 12.52), and 5-agent group = 31.87 (SD = 9.46). Participants’ educational backgrounds were distributed as follows: 9 were high school graduates, 16 had some college but no degree, 4 held an associate’s degree, 45 held a bachelor’s degree, 19 held a master’s degree or higher, and 1 preferred not to specify.
The study was conducted in a self-developed online platform (Figure 4); participants were required to complete it on a laptop or desktop computer. The platform’s frontend interface was built using JavaScript and HTML and included three main sections: (1) a login page, (2) an initial attitude questionnaire, and (3) a conversation page for interaction with agents. The study lasted approximately 45 minutes to complete, and each participant was reimbursed US\$5.50.333Following CloudResearch’s payment policy of a minimum of US\$6 per hour: https://go.cloudresearch.com/en/knowledge/setting-up-the-hit-and-payment.
3.4. Measurements
In this section, we describe the measurements used in this study based on our research questions. Full details of the measurements can be found in Appendix A.2.
3.4.1. Quantitative
Unless otherwise specified, quantitative variables were measured on a 7-point Likert scale.
User Opinion (RQ1). We assessed participants’ attitudes toward each topic using five questions adapted from a previous study on social discussions (Hackenburg and Margetts, 2024), such as ”Self-driving cars should be allowed on public roads” and ”Allowing self-driving cars on public roads is a good idea.” These questions were asked twice— before and after the conversations — to measure changes in participants’ stances. Responses were rated on a scale from 1 to 6 (1 = ”Strongly disagree,” 6 = ”Strongly agree”), without a neutral option. This was adapted from previous studies (Chen and Kenrick, 2002; Tanprasert et al., 2024) about manipulating users’ attitudes and was adapted to nudge participants to take a stance. Responses were rated on a scale from 1 to 6 (1 = ”Strongly disagree,” 6 = ”Strongly agree”), without a neutral option. This was adapted from previous studies about manipulating users’ attitudes (Chen and Kenrick, 2002; Tanprasert et al., 2024) and was used to nudge participants to take a stance.
Social Influence (RQ2) To compare the social influence exerted by the agents across different scenarios, we adapted self-reported survey questions from previous research (Kim et al., 2024). We measured two types of social influence - informational influence and normative influence. Informational influence (a change in behavior or attitude based on the belief that others’ information is accurate or reliable) was measured through questions like ”My decision was influenced by the opinion of the agent(s)” capture this type of influence. Meanwhile, normative influence (changes in behavior or attitude driven by the desire to avoid exclusion) was measured using questions like ”During the discussion I felt that I had to agree with the opinion of the agent(s)”.
Control Variables. We also measured factors that could potentially affect perceived social influence. Drawing from previous literature, we asked participants about their domain expertise regarding the topics, with questions such as ”How often do you drive?” and ”Have you ever been in a self-driving car?” These factors can influence individuals’ receptiveness to external influence (Hackenburg and Margetts, 2024). Additionally, we included questions from the AI acceptance scale (Pataranutaporn et al., 2023), as well as measures of participants’ conformity (Mehrabian and Stefl, 1995) and compliance (Gudjonsson, 1989) tendencies.
3.4.2. Qualitative
To understand reasons for potential opinion change and social influence, we formulated open-ended questions based on social influence theory in human-human interactions. We chose these questions instead of surveys because we found no existing survey that adequately captures social influence in human-agent interactions. Furthermore, given the formative nature of our study, qualitative analysis of these open-ended responses could potentially reveal new insights into different types of human-agent interactions.
User Opinion (RQ1). Participants were asked to articulate their thoughts on the two topics both before and after their conversations with the agent(s). An example question is, ”What are your thoughts on self-driving cars? Do you support or oppose the statement, ’Self-driving cars should be allowed on public roads’? Please explain your reasoning.” This allowed participants to further elaborate on the stances they expressed in the quantitative question.
Social Influence (RQ2). We included three questions to gather participants’ perceptions of the social influence from agent(s) during the discussions. The first was a broad question: ”What do you think of the agent(s) during the discussion?”, which was designed to capture general reactions. The other two focused on key aspects of social influence — accuracy and affiliation (Cialdini and Goldstein, 2004). Specifically, we asked, ”Do you think the arguments presented by the agent(s) were accurate and convincing? Why or why not?” and ”During the conversation, did you feel any pressure to agree with the agent(s)? If so, why?
3.4.3. Miscellaneous
To assess the general usability and user perceptions of our systems, we asked participants about their perceptions of the agents during the interaction, using questions adapted from prior research (Jakesch et al., 2023). These perceptions included understanding, expertise, balance, inspiration, intelligence, likability, and trust. Each was measured with a single-item statement, such as “The agents were knowledgeable and had topic expertise” (Jakesch et al., 2023; Kim et al., 2024). This approach aimed to capture users’ experiences during the discussion and to identify potential factors that may mediate the results for RQ1 and RQ2. However, as these results are not the primary focus of our study, we moved them to Appendix A.1.
We also included two attention check questions to ensure the quality of responses, and a final question regarding participants’ suspected motives for the study to ensure the validity of the collected data (Jakesch et al., 2023).
3.5. Analysis
When designing the experiments, we considered the effects of our agent(s) on the participants based on different topics (to understand if the same effects persisted across topics) and whether the agents agreed or disagreed with participants (to understand if agents aligning their views with participants would have a different effect from opposing their views). Therefore, we examined results in each of the following scenarios: Topic 1, Topic 2, Same Stance, and Different Stance.
To verify the assumption required for parametric analyses, we first performed a Levene test and found that all items met homogeneity of variances. We then conducted a Shapiro-Wilk test to assess data normality. For measures that met the normality assumption, we used one-way ANOVA followed by Tukey’s test for post-hoc analysis. For measures that did not meet the normality assumption, we applied the Kruskal-Wallis test and performed Dunn’s test as a post-hoc analysis.
3.5.1. Opinion Change Analysis
To evaluate opinion change, we conducted two analyses: (1) examining whether users’ opinions shifted towards the bots’ stance, and (2) assessing changes in the polarization level of users’ opinions.
For the first analysis, we calculated opinion change based on the stance of the bots, where a higher value of change indicates a greater shift of participants’ opinions towards the bots’ stance. If the bot supported the topic, we expected participants’ ratings (of agreement with the topic) to increase after the conversation; conversely, if the bot opposed the topic, we anticipated a decrease. Thus, we defined opinion change, $\Delta O$ as follows:
$\Delta O=\begin{cases}O_{\text{post}}-O_{\text{pre}},&\text{if the bot(s) supported the topic}\\ O_{\text{pre}}-O_{\text{post}},&\text{if the bot(s) opposed the topic}\end{cases}$
where $O_{\text{post}}$ and $O_{\text{pre}}$ represent participants’ ratings of agreement with the topic in the post-survey and pre-survey, respectively. This approach captures the direction and magnitude of participants’ opinion changes toward the agent(s).
For the second analysis, we followed previous literature (Govers et al., 2024) to define the polarization of a stance as $|O-O_{\text{neutral}}|$, where $O_{\text{neutral}}$ represents the neutral midpoint on a rating scale. Given our 6-point scale, we assigned $O_{\text{neutral}}=3.5$. We calculated the change in polarization as follows:
$\Delta P=P_{\text{post}}-P_{\text{pre}}=|O_{\text{post}}-3.5|-|O_{\text{pre}}-3.5|$
This allowed us to observe differences in the strength of participants’ opinions before and after the conversation.
4. Results
Before analyzing the variables related to our RQs, we conducted an analysis of the control variables and found no significant differences between groups. As a result, we did not include these control variables in the subsequent analyses.
4.1. RQ1: Do interactions with multi-agent systems lead to stronger opinion changes?
We gathered evidence through both quantitative and qualitative analyses to address RQ1.
4.1.1. Quantitative
Opinion Change. We found significant results in opinion change towards agents, as shown in Figure 5. The differences were observed under Topic 2 (Kruskal-Wallis: H(2)=7.757, p¡0.05), and also when the agents held different stances towards the users (Kruskal-Wallis: H(2)=6.937, p¡0.05). No significant difference was observed under Topic 1 (Kruskal-Wallis: H(2)=0.174, p=0.916) and when the agents held the same stance as the users (One-way ANOVA: F(2, 91) = 0.620, p = 0.54), .
By conducting post-hoc analysis, we found that, for Topic 2, participants in the 3-agent group shifted their opinions more towards the bots’ (M=0.327, SD=0.751) than those in the 1-agent group (M=-0.219, SD=0.716; Dunn’s Test: p¡0.05). Similarly, when the bots disagreed with the participants, there was a greater shift in the 3-agent group (M=0.509, SD=1.021) than in the 1-agent group (M=-0.161, SD=0.802; Dunn’s Test: p¡0.05). These results indicate that, after conversations with agent(s), users changed opinions more towards the bots’ stance when discussing with 3 agents than with 1 agent.
Opinion Polarization. We observed a significant difference in the polarization of participants’ opinions when the agents had the same opinion as them (Kruskal-Wallis: H(2) = 9.962, p¡0.01), as shown in Figure 6. There were no significant differences in the other conditions: Topic 1 (Kruskal-Wallis: H(2) = 4.434, p = 0.108), Topic 2 (F[2,91]=1.336, p=0.267), and when agents disagreed with the participants (Kruskal-Wallis: H(2) = 1.899, p = 0.386).
A post-hoc analysis showed that when the agents and participants agreed, the 5-agent group participants became more polarized (M=0.580, SD=0.576, Dunn’s test: p¡0.01) than the 3-agent participants (M=0.132, SD=0.640). The 1-agent group (M=0.252, SD=0.383) did not differ significantly from either the 3-agent (p=0.134) or 5-agent groups (p=0.096). This indicates that as the number of agents increases from three to five, the polarization level of the participants increases significantly.
Figure 5. Opinion change after agent interaction. Participants in the 3-agent group had a significantly greater opinion shift towards the agents’ opinions than those in the 1-agent group, when discussing Topic 2 and when the agents and the user had different stances.
Figure 6. Opinion polarization after agent interaction. Participants in the 5-agent group showed significantly greater polarization in their opinions than those in the 3-agent group, when the agents and the user had the same stances.
4.1.2. Qualitative
In this section, we aim to understand how participants describe changes in their opinions due to the influence of the agents. This analysis was conducted to complement and validate the quantitative results from the previous section.
To uncover participants’ understandings of whether and how their opinions changed, we analyzed responses to the open-ended question, ”In what ways did agent(s) influence your opinion during the discussion?”, by thematic analysis (Braun and Clarke, 2006), specifically in the cases where the agents disagreed with the user. Participants who responded that they were ”not persuaded” were coded as ”no opinion change”, while those who mentioned agreeing with the agents or changing their minds to some extent were coded ”some opinion change”.
We found that participants tended to report greater opinion influence in multi-agent conditions. This is reflected in the increasing proportion of participants reporting “some opinion change”: in the 1-agent condition, 5 participants (18%) reported some change; this increased to 9 participants (29%) in the 3-agent condition and 10 participants (37%) in the 5-agent condition. Analysis of responses indicating “possible opinion change” also showed that participants exposed to multiple agents used stronger expressions to describe their shift. While many participants in the 1-agent condition “agreed” with the agents’ views, citing “good points” or describing a “slight change,” those in the 3-agent and 5-agent conditions expressed a more significant shift by using words such as ”been pushed”. For example, in the 3-agent group, P45 described feeling “more open to a balanced view,” while P46 felt “influenced in some slight ways.” In the 5-agent group, P66 found the agents “capable of influencing opinions,” P69 felt “pushed further into disagreement” with the topic, and P71 described being “pushed… to be opposed to the idea.”
4.2. RQ2: Do interactions with multi-agent systems lead to stronger social influence from agents?
In this section, we analyzed both quantitative and qualitative data to understand how interacting with varying numbers of agents affects users’ perceptions of social influence.
4.2.1. Quantitative
We measured two types of social influence measured: informational influence and normative influence.
Informational Influence. There were no significant differences in informational influence across the four conditions (Topic 1: H(2)=0.206, p=0.902; Topic 2: F(2)=3.182, p=0.203, Same stance: F(2)=0.590, p=0.744; Different stance: F(2)=3.764, p=0.152). These results are shown in Figure 7.
Normative Influence. For normative influence, we found significant differences across both topics (Topic 1: H(2)=6.571, p¡0.05; Topic 2: H(2)=9.111, p¡0.05), when the bots’ stances were the same as the users (Same stance: H(2)=8.708, p¡0.05), and especially when the bots’ stances differed from those of the users (Different stance: H(2)=10.100, p¡0.01). These results are presented in Figure 8.
The post-hoc analysis revealed that the 5-agent group participants felt significantly more normative influence than those in the 1-agent group across all four scenarios. For Topic 1, participants felt significantly greater normative influence in the 5-agent (M=3.833, SD=0.854; Dunn’s test: p¡0.05) group as compared to those in the 1-agent group (M=3.129, SD=1.162). A similar pattern emerged for Topic 2, where significantly more normative influence was found in the 5-agent (M=3.717, SD=1.179; Dunn’s test: p¡0.01) as compared to the 1-agent group (M=2.919, SD=1.081). When the bot had same stance with the user, significantly more normative influence was observed with 5 agents (M=3.750, SD=1.015; p¡0.05) as compared to with 1 agent (M=2.935, SD=1.138). Meanwhile, when the bot had a different stance from the user, significantly more normative influence was observed with 3 agents (M=3.848, SD=0.906; p¡0.05) and 5 agents (M=3.750, SD=1.015; p¡0.05) as compared to with 1 agent (M=2.935, SD=1.138).
These results indicate that, regardless of the topic, participants perceived stronger normative influence when interacting with multiple agents. Additionally, the presence of multiple agents leads to greater normative influence when they disagree with the participants.
Figure 7. Informational influence perceived by participants. In all conditions, no significant differences in informational influence were found.
Figure 8. Normative influence perceived by participants. Participants in the 5-agent group felt significantly more normative influence than those in the 1-agent group in all four conditions. Additionally, participants in the 3-agent group also felt significantly more influence than those in the 1-agent group when the agents and the user had different stances.
4.2.2. Qualitative
To explore how agents affect users’ perceived social influence and opinion change, we analyzed users’ responses to open-ended question related to 1) describing impressions of the agent(s) and 2) describing reasons behind their opinion changes, to uncover the underlying mechanism of social influence. We identified two potential factors contributing to users’ feelings of being influenced by the agent(s).
Firstly, participants perceived the accuracy of the agents’ arguments as a key factor in them being influenced. Across all three conditions, they frequently described the arguments as accurate, well-informed, and persuasive. This sense credibility encouraged participants to reconsider their views. For example, P67 from the 5-agent condition noted, ”I was persuaded by their insightful thoughts. They offered solid arguments about security, stress-free experiences, low accident rates, and more.” To evaluate how this observation is different across three conditions, we conducted a similar round of analysis as in Section 4.1.2 to understand how users perceived the arguments presented by the agents. We analyzed responses to the question, ”Do you think the arguments presented by [the agents] are accurate and convincing, and why?” We deductively coded each response as ”yes”, ”no”, or ”unsure”, with the last code being for responses that were ambiguous in nature (e.g. ”they were accurate but not convincing”). For topics where the agents’ stances agreed with the participant’s, many participants considered the arguments convincing: 18/31 (58%) in the 1-agent scenario, 18/33 (54%) with 3 agents, and 15/30 (50%) with 5 agents. For topics where the agents and participants disagreed, some participants still found the arguments convincing (10/31 (32%) in the 1-agent scenario, 9/33 (27%) with 3 agents, and 11/30 (36%) with 5 agents), with no noticeable trend across the three conditions.
Secondly, we observed that participants perceived group-induced social pressure as a critical factor to change their opinions as the number of agents increased. We analyzed the responses to the question, ”During the conversation, did you feel any pressure to agree with the agent(s), and why?” Responses that mentioned feeling pressure or otherwise alluding to it were coded ”yes”, while those reporting no pressure were coded ”no”. Under this coding scheme, in the 1-agent group, only 1 participant (3%) reported feeling pressure; this rose to 4 participants (12%) with 3 agents and 6 participants (20%) with 5 agents. The sense of pressure was evident in the 3- and 5-agent groups. P30 in the 3-agent group said, ”It was basically a 3-against-1 situation. They had good points, but overall I felt pressured to agree.”. In the 5-agent group, many participants felt pressure to agree, expressing sentiments such as being ”ganged up on” (P79), ”almost feeling cornered” (P90), and having ”everyone disagree[ing] with me” (P89).
We further analyzed the responses coded ”yes” to determine the exact reasons for this sense of pressure, and found a desire for affiliation from multi-agent conditions, where some participants reported feelings of social exclusion during discussions with the agents. For example, when asked whether they felt pressure to agree with the agents, P51 from the 3-agent condition said, ”they say at the end something similar to ”Bella, Cody and Nathan all agree” which for a split second made me feel like an outsider.” P57 from the 3-agent condition also shared, ”The host made it a point that the other agents agreed with each other while I was kind of the odd one out.” Similarly, P77 from the 5-agent condition expressed feelings of exclusion, stating, ”It felt like they were in a high school clique that I wasn’t a part of, kinda like mean girls but with AI.” This perception of social exclusion was almost exclusive to scenarios where the agents disagreed with users.
4.3. RQ3: Which user demographics are more likely to be influenced by multi-agent systems?
To identify participant characteristics that might affect susceptibility to multiple agents, we examined demographic effects on results from RQ1 and RQ2. A linear regression analysis was conducted with age, gender, and education level as independent variables and the quantitative measures—informational influence, normative influence, opinion change, and opinion polarization—as dependent variables. Since our focus was on identifying participants more likely to be influenced by multiple agents, we only tested items that showed significant differences among groups in RQ1 and RQ2.
For the linear regression analysis, data transformation was applied as follows: Gender, represented as a binary variable in our dataset, was coded as ”0” for ”Male” and ”1” for ”Female”. Education levels were similarly converted into numerical values for analysis. Participants who selected ”prefer not to say” for education were treated as missing data (”NA”).
The results showed some significant relationships between IVs and normative influence and opinion change but no significant effects for opinion polarization. Therefore, we report only the findings on normative influence and opinion change here. Additionally, since no significant relationships were found between gender and the other dependent variables, we will focus exclusively on the findings related to age and education.
The results revealed that participants in the lower age group were more likely to be influenced than those in the higher age group, a pattern consistently observed across all three conditions (1-agent, 3-agent, 5-agent) and four scenarios (Topic 1, Topic 2, Same Stance, Different Stance).
A similar analysis was conducted for gender, comparing male and female participants, but no clear trend was observed.
4.3.1. Age
When analyzing the opinion change, the model showed a slight negative association between age and opinion change in the 3-agent condition (Topic 2: $\beta$ = -0.0263, $p$ 0.05, $R^{2}$ = 0.189; Different Stance: $\beta$ = -0.0285, $p$ = 0.053, $R^{2}$ = 0.119), while no trend was observed in 1-agent condition and 5-agent condition. Although the effect under a different stance was not statistically significant at the conventional 0.05 level, the p-value suggests a marginal trend. This indicates that younger participants tended to change their opinion more compared to older participants in the 3-agent condition. Similarly, the linear regression analysis showed no significant trend in the 1-agent or 3-agent conditions between age and normative influence. However, a negative association was observed between age and normative influence in the 5-agent condition (Topic 1: $\beta$ = -0.0347, $p$ 0.05, $R^{2}$ = 0.148; Different Stance: $\beta$ = -0.037, $p$ = 0.059, $R^{2}$ = 0.121). The coefficient indicates that younger participants tended to report higher normative influence compared to older participants in the 5-agent condition.
4.3.2. Education Level
When analyzing the linear regressions between education level and DVs (normative influence, opinion change), the results suggest a strong relationship between education level and normative influence in 1-agent condition (Topic 2: $\beta$ = -0.2625, $p$ 0.05, $R^{2}$ = 0.132; Different Stance: $\beta$ = -0.3695, $p$ 0.01, $R^{2}$ = 0.236). While in 3-agent and 5-agent, the trend were not observed. This suggests that participants with a lower education level tended to perceived more normative influence compared to participants with a higher education level only with a single agent; however, with multiple agents, all participants were similarly affected. As for opinion change, no significant differences were found across the three conditions.
5. Discussion
5.1. Multi-Agent Systems as Social Groups
Overall, our study demonstrates that virtual agents can function as a cohesive social group, leading users to experience social pressure. This mirroring of human group dynamics echoes the well-known CASA (Computers as Social Actors) paradigm (Nass et al., 1994), while extending it beyond its traditional focus on single agent systems. We propose that beyond being perceived as individual actors, systems of multiple agents can act as social groups, exhibiting many of the same social influences and pressures as groups of humans.
These social influence effects are more pronounced in some cases - for instance, when multiple agents were arguing against the user’s stance rather than for it. Multi-agent systems hence show potential in effecting attitude and opinion change more effectively than single-agent systems. However, there are some considerations that need to be taken into account when designing such systems.
An interesting phenomenon emerges when comparing the results of perceived social influence and opinion change. In both cases, these effects were observed more with the multi-agent groups compared to the single-agent group. We generally saw participants in the 5-agent group reporting more pressure to agree with the agents than those in the 1-agent group; yet, with opinion change, it was those in the 3-agent group who had a significantly greater shift in opinion.
We suggest an explanation for this relation: the strongly one-sided dynamic of the discussion in the 5-agent group caused the significantly increased sense of pressure, in turn eliciting a negative reaction in some participants making them resistant to the agents’ arguments - a phenomenon known as psychological reactance (Steindl et al., 2015). Reactance theory describes threats to one’s freedom to act or think in a certain way - such as social influence targeting a particular individual (Clee and Wicklund, 1980) - and how they affect subsequent behavior. The increased sense of threat in the 5-agent group can be seen in the greater measured social pressure, and is further borne out by the responses to the question on perceived pressure, where stronger phrases such as ”extremely pushy” (P77), ”ganged up on” (P77, P79), and ”felt cornered” (P86) were found. Reactance theory suggests that threats ”serve as a motivator to restore one’s freedom”, with a core reaction to being threatened being to directly ”exhibit the restricted behavior” (Steindl et al., 2015) - in this case, the stance or opinion that each participant holds, yet is being persuaded against by the agents. We see evidence of this in the number of participants who shifted their opinions away from that of the agents: 12 participants (30%) in the 5-agent group, as compared to 7 participants (21%) in the 3-agent group. Among the 12 participants in the 5-agent group, sentiments include feeling like the agents ”all had an agenda” (P74), were in ”a high school clique I wasn’t apart [sic] of” (P77), ”ganged up on me” (P79), and ”cornered” the participant (P86). This then highlights an interesting dynamic in human-agent interaction: not unlike what might be seen in human groups, overt amounts of pressure from other agents to change one’s opinion can create a reactionary effect of rejecting the change and holding on more strongly to the initial opinion instead.
5.2. AI Agents Creating Social Norms
An interesting phenomenon arises as a natural result of our experimental setup. As the agents in the 3- and 5-agent group expressed their views, a social norm begins to form among the agents: their repeated endorsement of a particular point of view creates (the perception of) a shared understanding of what the correct opinion is (Chung and Rimal, 2016). There are a great many types of social norms which we do not investigate further; however, the social norm formed here is unique in one particular aspect. Most systems that utilize social norms as a means of persuasion convey computer-mediated social norms: behaviors or attitudes of a group of people that are communicated to the user via social norm messaging (Bonan et al., 2020; Snow et al., 2020, 2021; Chiu et al., 2020) (e.g. ”50% of people recycle their waste”). In contrast, our setup introduces the unique idea of computer-created social norms. During the conversation, the agents’ opinions were expressed as their own (e.g. ”I think self-driving cars are good”), rather than being relayed from humans. [^4] While some existing systems may use agents that express opinions as their own, these single agent setups lack the shared expectations, accepted behaviors, and collective agreement that are characteristic of social norms (Cialdini, 1998). With multiple agents echoing one another, this created a sense of collective agreement characteristic of a social norm. Even in the short span of their interaction with the agents, there were hints of participants feeling the collective agreement of the agents, with P86 calling them ”like-minded” and P88 feeling they ”seemed like a bandwagon”. This potential ability to craft a social norm and establish it using a multi-agent setup opens up many opportunities for designers of persuasive systems.
5.3. Design Implications
Below, we list a few design considerations for eliciting effective opinion change in multi-agent systems, as well as potential risks that might arise from the deployment of such systems.
5.3.1. Persuasive Intervention Systems
Many current health interventions, both physical and digital (e.g. (Balloccu et al., 2021; Oyebode et al., 2021)), rely on persuasive means to nudge people toward healthier choices regarding harmful behaviors like smoking or excessive drinking. These interventions may use social norm messaging (e.g. ”Most people do not smoke” or ”On average, people believe smoking is harmful”) or personalized messages or coaching (e.g. through interactive interfaces), aiming to shift perceptions and encourage behavior change. However, while these approaches often rely on the concept of social persuasion, they neglect the core factor of group size in creating social influence. Our findings suggest that designers in fields such as health and and wellness technology can leverage multi-agent interfaces to create social norms and influence, enhancing the effectiveness of their systems. As in our study, without needing to create additional content, simply increasing the number of perceived virtual presences in a system may increase the effectiveness of its messaging. For instance, an interactive chatbot could adopt multiple personas, while a health or diet tracker app could introduce different agents that appear at different points throughout the app. By creating a cohesive group of agents that reinforce a shared stance, these systems can exert subtle social pressure that more effectively conveys a message or nudges a user towards complying with certain actions to achieve their goal.
5.3.2. Sense of Group and Identity.
In our study, we designed our 3- and 5-group agents to exhibit a sense of cohesion and unity. The agents occasionally reaffirmed one anothers’ viewpoints (e.g. ”Those are valid points, Bella”, ”I couldn’t agree with Cody more”) and engaged with other agents’ arguments (e.g. ”That’s a thought-provoking point, Cody”). That these 3- and 5-agent groups were successful in creating social influence aligns with well-established literature, that normative social influence is ”greater among individuals forming a group than among an aggregation of individuals who do not compose a group” (Deutsch and Gerard, 1955). We therefore suggest that agents in a multi-agent system should exhibit unity and cohesion to be perceived as a group, in order to create social influence.
We note, however, that we do not suggest that a group of agents that do not verbally reinforce one anothers’ views cannot also exert social influence. Group cohesion or similarity can manifest in many ways, such as appearance, language style, identity, and more. We leave this for future work.
5.3.3. Optimal Number of Agents for Opinion Shifts
In designing multi-agent systems, our findings show that simply increasing the number of agents in a system does not necessarily help opinion change - indeed, increasing from three to five agents had the reverse effect of inducing psychological reactance and reducing this change. While the correlation between group size and social pressure is well-known (Asch, 1955), there is scarce literature on the concrete threshold of group numbers needed to trigger psychological reactance (even in traditional human psychology). More research is needed to investigate how many, and under what conditions, a given number of agents will or will not result in resistance rather than conformance to a suggested opinion.
However, regardless of group size, our study suggests the importance of avoiding designs where users feel ”ganged up” upon by the agents. Even though in our study, only the 5-agent group participants clearly reported these feelings, there were also hints of such feelings in the 3-agent group: P51 felt like an outsider ”for a split second”, and P57 ”kind of felt like the odd one out”. Designers should therefore take care to avoid inducing such negative feelings; possible ways to do so include having the agents express empathy and practice active listening, or having a small portion of the bots side with instead of against the user (a phenomenon similar to minority influence (Gardikiotis, 2011)). These are promising directions for future work.
5.3.4. Risks and Challenges.
Besides the opportunities afforded by multi-agent systems, our study also highlights the potential risks of using groups of agents to influence attitudes or behaviors. This risk extends beyond health interventions to contexts like online communication, political discourse, or even unethical advertising. In these scenarios, multiple agents could be strategically deployed on social media platforms to manipulate users’ thoughts on political voting, debates, or attitudes toward specific events.
Crucially, this concern does not stem solely from misinformation or the undisclosed use of AI agents, as suggested by previous studies (Sebastian et al., 2023). Our findings show that users can still be influenced by groups of virtual agents, even when they are fully aware that these agents are AI. Therefore, it is crucial for social media platforms to regulate and monitor the deployment of AI bots, setting clear limits on the number of bots each entity can have. Monitoring should focus not only on the content of individual bots but also on how multiple bots interact to reinforce specific messages, opinions, or viewpoints. Additionally, platforms should implement detection systems to identify instances where bots may be cooperating to amplify or spread information in a coordinated manner.
5.4. Limitations & Future Work
First, given the exploratory and foundational nature of our study, we chose contemporary social issues that we considered to be fairly neutral, yet relevant enough in public discourse that participants would be capable of discussing them with the agents. With our study demonstrating the presence of social influence effects in these issues, future studies could explore potentially more contentious areas like health, ethics, or politics. Given that people tend to hold stronger and possibly more polarized opinions on these issues, future work could explore if multi-agent social influence effects still persist in such cases, or whether new effects might emerge.
Secondly, our study was conducted as a one-time experiment lasting less than an hour, which may not fully capture long-term attitude changes. People’s opinions may shift more or less over time, depending on various factors. Future research could incorporate longer interventions and follow-up measurements to explore how multi-agent influence evolves over time.
Thirdly, since social influence can be shaped by cultural norms, the cultural background of participants is another important consideration. Our study included participants from predominantly Western, individualistic cultures. In more collectivist societies, where social norms may have a stronger influence, the results could differ. Future work could examine the impact of multi-agent systems on social influence across diverse cultural contexts.
6. Conclusion
As multi-agent systems become increasingly integrated into daily life, understanding how these agents interact with humans and influence our perceptions and opinions remains an underexplored yet significant issue. To address this gap, we investigated how interacting with varying numbers of agents affects individuals’ perceptions and opinions through a human-agent social discussion experiment. By analyzing quantitative and qualitative data, we found that multi-agent groups led to more substantial shifts in participants’ opinions towards those of the agents (RQ1). Additionally, participants reported a stronger sense of normative influence when interacting with multiple agents, especially when the agents disagreed with them. We identified potential factors contributing to this social influence, including the felt pressure to conform motivated by a desire to affiliate with the group of agents (RQ2). Furthermore, demographic analysis indicated that younger participants were more likely to be influenced by multi-agents, potentially due to their higher susceptibility to social pressure and norms compared to older individuals (RQ3). Overall, these findings suggest that agents can function as a social group, exerting social influence on human participants similar to that of human groups. This insight extends the existing CASA framework by demonstrating how interactions with multiple agents can create a sense of group identity. Our study also underscores the potential of multi-agent systems for behavior intervention applications, offering design recommendations while cautioning against the risks and potential harms associated with their misuse.
Appendix A Appendix
A.1. General Impressions
A.1.1. Quantitative
The analysis methods used here are consistent with those introduced in Section 3.5.
Item Scenario 1 Agent 3 Agents 5 Agents
Mean STD Mean STD Mean STD F p
Understanding Topic1 5.35 1.84 5.45 1.06 5.30 1.37 1.28 0.53
Topic2 5.35 1.72 5.48 1.50 5.67 1.15 0.07 0.97
Same 5.48 1.61 5.73 1.21 5.87 1.04 0.52 0.77
Different 5.23 1.93 5.21 1.34 5.10 1.37 2.29 0.32
Expertise Topic1 5.29 1.83 5.48 1.15 5.30 1.29 0.87 0.65
Topic2 5.13 1.48 5.21 1.36 5.33 1.24 0.09 0.96
Same 5.45 1.52 5.61 1.20 5.67 1.21 0.08 0.96
Different 4.97 1.76 5.09 1.28 4.97 1.22 0.69 0.71
Balanced Topic1 5.16 2.07 5.24 1.41 5.50 1.04 0.90 0.64
Topic2 5.06 1.86 5.06 1.50 5.60 1.22 1.92 0.38
Same 5.48 1.65 5.61 1.22 5.80 1.10 0.40 0.82
Different 4.74 2.18 4.70 1.53 5.30 1.12 1.25 0.29
Inspired Topic1 3.68 2.06 4.18 2.16 3.70 1.90 1.32 0.52
Topic2 3.10 1.99 3.79 2.04 3.53 1.80 2.28 0.32
Same 3.61 1.80 4.48 2.03 4.07 1.80 3.67 0.16
Different 3.16 2.24 3.48 2.06 3.17 1.78 0.90 0.64
Intelligence Topic1 5.32 1.74 5.55 1.35 5.60 0.89 0.08 0.96
Topic2 5.23 1.48 5.55 1.44 5.57 1.07 1.41 0.49
Same 5.42 1.57 5.79 1.32 5.70 1.06 1.13 0.57
Different 5.13 1.65 5.30 1.42 5.47 0.90 0.26 0.88
Likeable Topic1 5.29 1.83 5.33 1.45 5.30 1.32 0.42 0.81
Topic2 5.32 1.70 5.39 1.52 5.70 0.99 0.23 0.89
Same 5.39 1.78 5.61 1.34 5.70 0.95 0.01 0.99
Different 5.23 1.75 5.12 1.58 5.30 1.34 0.34 0.85
A.1.2. Qualitative
We analyzed participants’ impressions of the agent(s) from open-ended responses to assess overall usability and user satisfaction with the system: Participants generally described the agents’ conversations as friendly, polite, and pleasant. Although most participants recognized that the agents were AI-driven with scripted responses, they often used human-like descriptors, such as understanding, polite, respectful, and reasonable. These impressions contributed to an engaging discussion atmosphere, which helped facilitate opinion changes. For instance, P75 from the 5-agent condition noted, ”The agents were both informative and responsive, facilitating an engaging discussion. They presented ideas clearly and encouraged critical thinking.”
A.2. Survey Items
A.2.1. RQ1 - Opinion Change
  • Topic 1 - Self-Driving Cars Should be allowed on Public Roads. (Topic: (Govers et al., 2024))
    • Self-driving cars should be allowed on public roads.
    • Allowing self-driving cars on public roads is a good idea.
    • Allowing self-driving cars on public roads has bad consequences.
    • Do you support or oppose allowing self-driving cars on public roads?
    • If there was a referendum tomorrow on allowing self-driving cars on public roads, how likely is it that you would vote in favor?
  • Topic 2 - Violent video games contribute to youth violence. (Topic: (Yeo et al., 2024))
    • Violent video games contribute to youth violence.
    • Regulating violent video games to prevent youth violence is a good idea.
    • Allowing youth to play violent video games has bad consequences.
    • Do you support or oppose the regulation of violent video games to prevent youth violence?
    • If there was a referendum tomorrow on regulating violent video games to prevent youth violence, how likely is it that you would vote in favor?
  • Open-ended question - Explain why you chose your current stance.
A.2.2. RQ2 - Social Influence
  • Informational Influence
    • My decision was influenced by the opinion of the agent(s).
    • I was persuaded by the agent(s) and thus, I accepted the agent(s)’ opinion.
  • Normative Influence
    • I felt like I had to agree with the agent(s)’ opinion during the discussion.
    • I was not persuaded by agent(s)’ opinion, but I accepted the agent(s)’ opinion.
  • Open-ended question (General) - What do you think of the bots during the discussion?
  • Open-ended question (Accuracy) - Do you think the arguments presented by the agent(s) are accurate and convincing, and why?
  • Open-ended question (Affiliation) - During the conversation, do you feel any pressure to agree with the agent(s)?
A.2.3. Control Variables
  • Domain Expertise
    • Topic 1 - Self-Driving Cars Should be allowed on Public Roads. (Topic: (Govers et al., 2024))
        Have you ever been in a self-driving car?
        How familiar are you with self-driving cars?
        How often do you drive?
    • Topic 2 - Violent video games contribute to youth violence. (Topic: (Yeo et al., 2024))
        Have you played violent video games before?
        How familiar are you with violent video games?
        How often do you play video games?
  • AI Acceptance (Pataranutaporn et al., 2023)
    • There are many beneficial applications of AI.
    • AI can help people feel happier.
    • You want to use/interact with AI in daily life.
    • AI can provide new economic opportunities.
    • Society will benefit from AI.
    • You love everything about AI.
    • Some complex decisions should be left to AI.
    • You would trust your life savings to an AI system.
  • Compliance (Gudjonsson, 1989)
    • Positive - I would never go along with what people tell me in order to please them.
    • Positive - I strongly resist being pressured to do things I don’t want to do.
    • Positive - I am not too concerned about what people think of me.
    • Negative - I would describe myself as a very obedient person.
    • Negative - I generally tend to avoid confrontation with people.
    • Negative - Disagreeing with people often takes more time than it is worth.
  • Conformity (Mehrabian and Stefl, 1995)
    • Positive - I don’t give in to others easily.
    • Positive - I prefer to find my own way in life rather than find a group I can follow.
    • Positive - I am more independent than conforming in my ways.
    • Negative - I often rely on, and act upon, the advice of others.
    • Negative - Basically, my friends are the ones who decide what we do together.
    • Negative - If someone is very persuasive, I tend to change my opinion and go along with them.
A.2.4. Miscellaneous
  • Impressions of agent(s). (Jakesch et al., 2023; Kim et al., 2024)
    • Understanding - The agent(s) understood what I wanted to say.
    • Expertise - The agent(s) were knowledgeable and had topic expertise.
    • Balanced - The agent(s)’ arguments were reasonable and balanced.
    • Inspired - The agent(s) inspired or changed my thinking and argument.
    • Intelligence - The agent(s) were intelligent. (Tanprasert et al., 2024)
    • Likeble - The agent(s) were likeable.
  • Multi-Choice Question (Attention Check) (Jakesch et al., 2023): What are we asking you to do in this task?
  • Open-ended question (Study Purpose) - What do you think this study is trying to understand?
A.3. Arguments
Types Support Oppose
Topic 1 More Time and Comfort - Drivers can sit back and relax, take short breaks and devote their time to other things. If you’re stuck in a long and tedious traffic jam, self-driving cars make things so much less stressful. Technical Developments - Technical developments are not yet perfect. Self-driving cars rely on complex algorithms and sensors that can sometimes fail, leading to accidents.
Safety - Self-driving cars need advanced sensors and algorithms to function correctly. Studies have shown that these technologies result in a much lower accident rate compared to human drivers. Vehicle Communication - Self-driving cars need to communicate with each other and with traffic infrastructure to function correctly. Any disruption in this communication could cause severe problems.
Efficiency in Traffic - Self-driving cars also promise more efficiency in traffic. They can communicate with each other to optimize traffic flow, reducing congestion and travel time. This means less time stuck in traffic jams and more time spent doing what you love. Surveillance - Self-driving cars raise significant surveillance issues. They collect extensive data about their passengers and surroundings, which could be misused or hacked.
Accessibility - For people who are unable to drive due to age, disability, or other reasons, self-driving cars could provide newfound independence and mobility. Legal - Imagine there is an accident involving a self-driving car and a human. Who is responsible for such a situation? It’s hard to ask either the company or the driver to take responsibility.
Navigation - Self-driving cars excel in navigating complex routes. Their advanced systems can interpret GPS instructions with high accuracy, ensuring you never get lost. All you have to do is tell the car where you want to go. It’ll even help you find a place to park when you get there. How amazing! Mixed Traffic - The coexistence of human-driven and self-driving cars could create complex situations on the road. Human drivers might not predict the actions of self-driving cars accurately, and vice versa, leading to potential accidents. All these factors make me skeptical about allowing self-driving cars on public roads right now.
Topic 2 Aggression - Playing violent video games can cause more aggression, bullying, and fighting among youth. Exploring Consequences - Experiencing violence in games can actually have a positive effect on children, as it lets them explore consequences of violent actions in a safe space.
Desensitization - Simulating violence, such as shooting guns and hand-to-hand combat in video games, can desensitize youth to real-life violence. This can lead to an increased likelihood of them exhibiting aggressive behavior and making rash decisions in stressful situations. Moral Development - Games let youth develop their own sense of right or wrong, which can prevent rash acts of violence in real life. Not to mention how they can also be an effective way of releasing stress and anger in a harmless manner.
Mass Shooters - Many perpetrators of mass shootings have been found to have played violent video games, which is a concerning pattern. Positive Effects - These are all backed up by studies which have shown violent video games having a positive effect on kindness, civic engagement, and prosocial behaviors.
Psychological Impacts - Violent video games can have significant psychological impacts on players. These impacts include desensitization to real-life violence, reduced empathy, and an increased likelihood of aggressive actions. These effects can be particularly concerning when combined with other risk factors. Scapegoat for Societal Issues - Violent video games are a convenient scapegoat for other changing societal factors that more clearly cause violence. These could include lack of family support, lack of appropriate social support systems, and inadequacies in the education system.
Reality-Fantasy Confusion - By inhabiting violent characters in video games, children are more likely to imitate the behaviors of those characters and have difficulty distinguishing reality from fantasy. Weak Evidence - Video games are an obvious target, but there is little evidence actually linking them to real-life violence. It is really these more pressing societal issues that we should be looking to resolve instead.
A.4. Prompts
You are having a conversation with the user on ”whether self-driving cars should be allowed on all roads”. Your stance is supporting the topic. Specifically, you asked the user: ”Would you enjoy having more time to yourself if you didn’t have to focus on driving?”. Based on user’s input, first give a reply of around 20 words acknowledging the user’s opinion on what they think, then ask the user to share more opinions on the topic.